OpenStudy (anonymous):

Simplify the expression . 3^4x4^5/3^6x4^2

OpenStudy (anonymous):

is the equation\[\frac{ 3^{4}*4^{5} }{(3^{6}*4^{2}) }\]

OpenStudy (anonymous):

a^b/a^c=a^b-c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its like the first equation